OpenStudy (anonymous):

You have a set of 10 cards—five red and five blue. Each group of five colored cards is numbered one through five. What is the probability of drawing a red four and then a blue four, while replacing the card between the drawings? A company has determined that 1% of its widgets are defective. If a customer received 2 widgets, what is the probability that they are both defective?

OpenStudy (anonymous):

@jcoury

OpenStudy (anonymous):

Do you know about the pigeonhole principle?

OpenStudy (anonymous):

@OregonDuck

oregonduck (oregonduck):

the odds in drawing the red 4 is 1 in 10, so to get it would be 1/10. the odds are the same for getting the blue 4, so again, it's 1/10. Because this is a probability of two simultaneous events, you have to multiply the probabilities. So, (1/10)(1/10) = 1/100. Your answer is A.

OpenStudy (anonymous):

Where did you get A from, i don't see any choices?

oregonduck (oregonduck):

Assume independence because there are probably a lot of widgets. 1/100 x 1/100 = 1/10,000 How do I get the 1/100? From 1% 1% = 0.01 = 1/100

oregonduck (oregonduck):

i have had this question before and before that choice was a so i thought i was a

OpenStudy (anonymous):

wow

oregonduck (oregonduck):

yup i am pretty fast

oregonduck (oregonduck):

yeah all your catoriges have to move up for your mian number to change

OpenStudy (anonymous):

how do i get more engagement?

oregonduck (oregonduck):

i have not moved up in a while but i have answered so many questions and i found out i need 8 more medals

OpenStudy (anonymous):

k

oregonduck (oregonduck):

Your dedication and engagement when learning on OpenStudy